OpenStudy (redstealth29):

1. Kim's age is twice that of her sister. When you add Kim's age to her sister's age, you get 36. How old is each sister?

OpenStudy (redstealth29):

I'm sorry if this is not really my level, but I hate algebra and I hate these word questions.

jhonyy9 (jhonyy9):

Kim"s age let 2x and her sister x and you know that 2x+x=36 3x=36 x=36/3 =12 Kim"s age will be 2*12 = 24 her sister = 12 bc. 24+12=36 hope helped

OpenStudy (redstealth29):

Thank you thank you thank you! I was so stuck on this problem...:P